# Distributed Solving of Linear Quadratic Optimal Controller with Terminal State Constraint

### v3

@@ Abstract
This paper is concerned with the linear quadratic (LQ) optimal control of continuous-time system with terminal state constraint. In particular, multiple agents exist in the system which can only access partial information of the matrix parameters. This makes the classical solving method based on Riccati equation with global information suffering. The main contribution is to present a distributed algorithm to derive the optimal controller which is consisting of the distributed iterations for the Riccati equation, a backward differential equation driven by the optimal Lagrange multiplier and the optimal state. The effectiveness of the proposed algorithm is verified by two numerical examples.
keywords:
@@ 1 Introduction
Distributed control refers to situations where different agents access different information, share their information through network topology, and make decisions by using own and neighbors’ information. One of the most popular problems is consensus control which has been well studied [25, 26]. For the optimal control problems, [27] studied the distributed optimal control for discrete-time systems. [28] discussed the distributed consensus optimal control strategy for multi-agent systems with a general directed graph. [29] investigated a globally optimal distributed formation control algorithm. [30] considered the online distributed convex optimization problem with local dynamic loss function sums under unbalanced graphs. A distributed LQG optimal control problem was investigated in [31], where each controller has its own input matrix.
@@
In this paper, we will study the LQ optimal control problem with terminal state constraint where there exist multiple agents and each agent accesses partial information of the matrix parameters. The main contribution is to present a distributed algorithm to derive the optimal controller. Specifically, the distributed algorithm is consisting of the distributed iterations for the Riccati equation, a backward differential equation driven by the optimal Lagrange multiplier and the optimal state.
The remainder of the paper is organized as follows. Section II describes the studied problem. The explicit and the numerical solutions in centralized way are provided in Section III. Section IV gives the distributed algorithm for the optimal controller. Numerical example is given in Section V. The conclusion is outlined in the last Section. Detailed proofs of the results are given in Appendix.
2 Problem formulation
@@
where $x\in R^{n}$ is the state, $u\in R^{m}$ is the control input. $A$ and $B$ are constant matrices with compatible dimensions. $t\in[0,T]$ is the continuous-time variable. The initial state is prescribed as $x(0)=x_{0}$. The terminal state is predetermined as $x(T)=x_{T}$.
@@
Assumption 4.
There exists matrices $M_{i},i=1,\ldots,N$ such that $BR^{-1}B^{\prime}=\frac{1}{N}\sum_{i=1}^{N}M_{i}^{\prime}M_{i}$, where $M_{i}$ is accessible for agent $i,i=1,\ldots,N$.
@@
Problem.  The aim of the paper is to design a distributed algorithm for the optimal controller which minimizes the cost function (2) subject to system (1) and ensures the state to satisfy $x(0)=x_{0}$ and $x(T)=x_{T}$ for each agent $i$ under Assumptions 1-4.
Figure 1. The structural diagram of the Problem.
Remark 1.
The structural diagram of the studied Problem is shown in Figure 1, where agent $i,i=1,2,\ldots,N$ can only access its own $A_{i}$, $B_{i}$ and communicate with its neighbors. The problem has important applications in the fields of urban intelligent surveillance, disaster emergency response, and collaborative search and rescue. For example, in the distributed control problem of heterogeneous multi-agent systems (HMASs) consisting of unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs) [32], the dynamics of the $i$ agent can be described by
$\dot{x}_{i}(t)=\mathcal{A}_{i}x_{i}(t)+\mathcal{B}_{i}u_{i}(t),i=1,..,N.$ (3)
where $x_{i}(t)\in R^{n}$, $u_{i}(t)\in R^{m}$. By defining $x(t)=[x_{1}(t)^{\prime}~{}\cdots~{}x_{N}(t)^{\prime}]^{\prime}$, $u(t)=[u_{1}(t)^{\prime}~{}\cdots~{}u_{N}(t)^{\prime}]^{\prime}$, the system (3) can be expressed in the global form of system (1) with $A=\frac{1}{N}\sum_{i=1}^{N}A_{i}$, $B=\frac{1}{N}\sum_{i=1}^{N}B_{i}$, where
$A_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&\mathcal{A}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],B_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&\mathcal{B}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right].$ (12)
It is obvious that each agent makes a decision based on its local information $A_{i}$, $B_{i}$, which aligns with the problem studied in this paper.
3 Preliminaries
3.1 Explicit solution by using all the information
In this part, we present the optimal solution to the Problem by assuming all the information of matrix parameters are available for the agents. Noting that the terminal state is predetermined, thus it is necessary to ensure that the system can be driven from the initial state to the terminal state. To this end, we firstly present the definition of the reachability of system (1) as below.
Definition 1.
A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
@@ 3.1 Explicit solution by using all the information
$\dot{\beta}(t)=-[A^{\prime}-P(t)BR^{-1}B^{\prime}]\beta(t),~{}\beta(T)=\lambda^{*}$ (17)
@@
Proof. The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
3.2 Numerical solution by using all the information
@@
(2)$\lim_{n\rightarrow\infty}P^{n}(t)=P(t)$, where $P(t)$ satisfies (13).
Proof. The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
@@
Proof. The result follows directly from Lemma 3.1. $\blacksquare$
From Lemma 3.1 and Corollary 3.2, we have obtained the numerical solution to (13) and (14). Accordingly, the optimal solution (16) can be numerically obtained, and then the detailed algorithm is given in Algorithm 1.
@@
Input: Let the initial matrices be $P^{n}(T)=0,n=1,2,3,\ldots$, $P^{0}(t)=0,t\in[0,T]$, and select a sufficient small constant $\varepsilon>0$.
@@ 4. Solve the equation
$\dot{\beta}_{c}(t)=-Z^{n}(t)\beta_{c}(t),\beta_{c}(T)=\lambda_{c}^{*},$
@@ 5. Solve the equation
Proof. From Lemma 3.1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$. This further implies that there exist sufficient small constants $\varepsilon_{1}>0,\varepsilon_{2}>0$ such that
@@
$\|\beta_{c}(t)-\beta(t)\|\leq\varepsilon_{2},$
@@ 4 Main results
where $n=1,2,\cdots$, $k=1,2,\cdots$, $P_{i,0}^{n}(t)=0$, $P_{i.k}^{n}(T)=0$,
@@
while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}\leq\infty$, $\gamma>0$, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
@@
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|V_{i,k}^{n}(t)-V(t)\|$ $=$ $0$ (37)
Proof. The detailed proof is given in Appendix A. $\blacksquare$
@@
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|\Psi_{i,k}^{n}(t)-\Psi(t)\|$ $=$ $0$ (40)
Proof. The result follows directly from Lemma 4.1. $\blacksquare$
@@
whose convergence is given below.
Lemma 4.3.
@@
Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
@@
$-\lambda_{i,k-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}].$
where
[…]
and show the convergence in the following result.
Lemma 4.4.
@@
Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
@@
$\dot{\beta}_{i}^{\infty}(t)$ $=$ $-Z^{\infty}_{i}(t)\beta_{i}^{\infty}(t),$ (52)
@@
×$\times{W_{i}^{\infty}}^{\prime}W_{i}^{\infty}\beta_{i}^{\infty}(s)ds-x_{i,k-1}^{*}(t)]$
@@
and derive the following convergence result.
Lemma 4.5.
@@
Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
From Lemmas 4.1-4.5 and Corollary 4.2, we have established the distributed solutions to (13) and (14). This enables the distributed design of the optimal controller (16) as shown in Algorithm 2.
Theorem 3.
$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (56)
$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (57)
Proof. From Lemmas 3.1-4.1 and Corollary 4.2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$. Together with Lemmas 4.3-4.4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4},$
Accordingly, by using Lemma 4.5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (56) and (57). This completes the proof. $\blacksquare$
In the following, we generalize the above results to the standard LQ problem without terminal state constraint, i.e. minimizing (2) s. t. (1) with $x(0)=x_{0}$.
In this case, the optimal controller $u^{\diamond}(t)$ is
$u^{\diamond}(t)=-R^{-1}B^{\prime}P(t)x^{\diamond}(t),$ (58)
where $P(t)$ is the solution of equation (13), $x^{\diamond}(t)$ is the solution of the following equation:
$\dot{x}^{\diamond}(t)$ $=$ $[A-BR^{-1}B^{\prime}P(t)]x^{\diamond}(t),~{}x^{\diamond}(0)=x_{0}.$ (59)
@@
Input: Let the initial matrices be $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$, $P_{i,k}^{0}(t)=0$, $W_{i,0}=0$, $\lambda_{i,0}^{*}=0$, and $x_{i,0}^{*}(t)=0$ and select a sufficient small constant $\varsigma>0$.
@@
    $\dot{\beta}_{i}(t)=-Z_{i,k}^{n}(t)\beta_{i}(t),~{}\beta_{i}(T)=\lambda_{i,q}^{*},$
@@
    ×$\times W_{i,\varpi}^{\prime}W_{i,\varpi}\beta_{i}(s)ds-x_{i,w-1}^{*}(t)]$

### v4

@@ Abstract
This paper is concerned with the linear quadratic (LQ) optimal control of continuous-time system with terminal state constraint. In particular, multiple agents exist in the system which can only access partial information of the matrix parameters. This makes the classical solving method based on Riccati equation with global information suffering. The main contribution is to present a distributed algorithm to derive the optimal controller which is consisting of the distributed iterations for the Riccati equation, a backward differential equation driven by the optimal Lagrange multiplier and the optimal state. Furthermore, the proposed distributed iteration method is extended to solve the consensus control problem for heterogeneous multi-agent systems, achieving the globally optimal performance of the system. The effectiveness of the proposed algorithm is verified by two numerical examples, where the performance index under the proposed distributed controller is smaller than that under the commonly used consensus control.
Index Terms:
@@ I Introduction
Distributed control refers to situations where different agents access different information, share their information through network topology, and make decisions by using own and neighbors’ information. One of the most popular problems is consensus control which has been well studied [25, 26]. For the optimal control problems, [27] studied the distributed optimal control for discrete-time systems. [28] discussed the distributed consensus optimal control strategy for multi-agent systems with a general directed graph. [29] studied the design of suboptimal distributed control for the identical decoupled dynamical systems. [30] considered the online distributed convex optimization problem with local dynamic loss function sums under unbalanced graphs. A distributed LQG optimal control problem was investigated in [31], where each controller has its own input matrix.
@@
In this paper, we will study the LQ optimal control problem with terminal state constraint where there exist multiple agents and each agent accesses partial information of the matrix parameters. The main contribution is to present a distributed algorithm to derive the optimal controller. Specifically, the distributed algorithm is consisting of the distributed iterations for the Riccati equation, a backward differential equation driven by the optimal Lagrange multiplier and the optimal state. Furthermore, we solve the optimal consensus control problem using the proposed distributed iteration method. Numerical examples demonstrate the effectiveness of the proposed algorithms, where the performance index under the proposed distributed controller is smaller than that under the commonly used consensus control.
The remainder of the paper is organized as follows. Section II describes the studied problem. The explicit and the numerical solutions in centralized way are provided in Section III. Section IV gives the distributed algorithm for the optimal controller. Section V gives the distributed solution to the optimal consensus control problem. Numerical examples are given in Section VI. The conclusion is outlined in the last Section. Detailed proofs of the results are given in Appendix.
The following notations will be used throughout this paper: $R^{n}$ denotes the set of $n$-dimensional vectors; $x^{\prime}$ denotes the transpose of $x$. II is identity matrices with compatible dimensions.
II Problem formulation
@@
where $x(t)\in R^{n}$ is the state, $u(t)\in R^{m}$ is the control input. $A$ and $B$ are constant matrices with compatible dimensions. $t\in[0,T]$ is the continuous-time variable. The initial state is prescribed as $x(0)=x_{0}$. The terminal state is predetermined as $x(T)=x_{T}$.
@@
Assumption 4
There exists matrices $M_{i},i=1,\ldots,N$ such that $BR^{-1}B^{\prime}=\frac{1}{N}\sum_{i=1}^{N}M_{i}M_{i}^{\prime}$, where $M_{i}$ is accessible for agent $i,i=1,\ldots,N$.
@@
Problem 1. The aim of the paper is to design a distributed algorithm for the optimal controller which minimizes the cost function (2) subject to system (1) and ensures the state to satisfy $x(0)=x_{0}$ and $x(T)=x_{T}$ for each agent $i$ under Assumptions 1-4.
Remark 1
It is noted that Assumptions 1-4 are meaningful and have practical applications in network systems [31] and multi-agent systems, where the application of multi-agent systems is studied in Section V. In addition, we also note that although the system matrix was assumed to have a decomposable form under Assumptions 1-4, solving Problem 1 cannot be simply decomposed into the sum of $N$ subproblems, for example, in distributed optimal control problem where Assumptions 1-4 hold naturally, the derivation of the optimal solution is NP-hard [29].
III Preliminaries
III-A Explicit solution by using all the information
In this part, we present the optimal solution to the Problem 1 by assuming all the information of matrix parameters are available for the agents. Noting that the terminal state is predetermined, thus it is necessary to ensure that the system can be driven from the initial state to the terminal state. To this end, we firstly present the definition of the reachability of system (1) as below.
Definition 1: A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
@@ III-A Explicit solution by using all the information
$\dot{\beta}(t)=-[A^{\prime}-P(t)BR^{-1}B^{\prime}]\beta(t),~\beta(T)=\lambda^{*},$ (7)
@@
Proof: The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
III-B Numerical solution by using all the information
@@
[…]
Proof: The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
@@
Proof: The result follows directly from Lemma 1. $\blacksquare$
From Lemma 1 and Corollary 1, we have obtained the numerical solution to (3) and (4). Accordingly, the optimal solution (6) can be numerically obtained, and then the detailed algorithm is given in Algorithm 1.
@@
Input: Let the initial matrices be $P^{n}(T)=0,n=1,2,3,\ldots$, $P^{0}(t)=0,t\in[0,T]$, and select a sufficiently small constant $\varepsilon>0$.
@@ 4. Solve the equation
$\dot{\beta}_{c}(t)=-Z^{n}(t)^{\prime}\beta_{c}(t),\beta_{c}(T)=\lambda_{c}^{*},$
@@ 5. Solve the equation
Proof: From Lemma 1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$. This further implies that there exist sufficient small constants $\varepsilon_{1}>0,\varepsilon_{2}>0$ such that
@@
$\|\beta_{c}(t)-\beta(t)\|\leq\varepsilon_{2}.$
@@ IV Main results
where $n=1,2,\cdots$, $k=1,2,\cdots$, $P_{i,k}^{0}(t)=0$, $P_{i,0}^{n}(t)=0$, $P_{i,k}^{n}(T)=0$,
@@
while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$, $\gamma>0$ is selected such that the matrix $I_{N}-\frac{1}{\gamma}\mathcal{L}-\frac{1}{N}1_{N}1_{N}^{\prime}$ is Hurwitz, ensuring the convergence of the distributed iterations, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
@@
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|V_{i,k}^{n}(t)-V(t)\|$ $=$ $0.$ (27)
Proof: The detailed proof is given in Appendix A. $\blacksquare$
@@
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|\Psi_{i,k}^{n}(t)-\Psi(t)\|$ $=$ $0.$ (30)
Proof: The result follows directly from Lemma 2. $\blacksquare$
@@
with $W_{i,0}=0$, whose convergence is given below.
Lemma 3
@@
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
@@
$-\lambda_{i,k-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}],$
with $\lambda_{i,0}^{*}=0$ and
$\varrho_{i}^{\infty}=\int_{0}^{T}\Phi^{\infty}_{i}(s){W_{i}^{\infty}}^{\prime}W_{i}^{\infty}\Psi^{\infty}_{i}(s)ds.$ (40)
The convergence of this iteration in the following result.
Lemma 4
@@
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
@@
$\dot{\beta}_{i}^{\infty}(t)$ $=$ $-Z^{\infty}_{i}(t)^{\prime}\beta_{i}^{\infty}(t),$ (42)
@@
$\times{W_{i}^{\infty}}^{\prime}W_{i}^{\infty}\beta_{i}^{\infty}(s)ds-x_{i,k-1}^{*}(t)]$
@@
with $x_{i,0}^{*}(t)=0$ and derive the following convergence result.
Lemma 5
@@
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
From Lemmas 2-5 and Corollary 2, we have established the distributed solutions to (3) and (4). This enables the distributed design of the optimal controller (6) as shown in Algorithm 2.
@@
Input: Let the initial matrices be $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$, $P_{i,k}^{0}(t)=0$, $W_{i,0}=0$, $\lambda_{i,0}^{*}=0$, and $x_{i,0}^{*}(t)=0$ and select a sufficiently small constant $\varsigma>0$.
@@
    $\dot{\beta}_{i}(t)=-Z_{i,k}^{n}(t)^{\prime}\beta_{i}(t),~\beta_{i}(T)=\lambda_{i,q}^{*},$
@@
    $\times W_{i,\varpi}^{\prime}W_{i,\varpi}\beta_{i}(s)ds-x_{i,w-1}^{*}(t)]$